# Efficient Large Language Models: A Survey

### v2

@@ 2.2. Efficient Pre-Training
Optimization Strategies. Popular LLMs such as GPT-3 (Brown et al., 2020), OPT (Zhang et al., 2022d), BLOOM (Scao et al., 2022), and Chinchilla (Hoffmann et al., 2022) are predominately pre-trained using Adam (Kingma and Ba, 2017) or AdamW (Loshchilov and Hutter, 2017) as optimizers. However, both Adam and AdamW have a huge demand on memory and are computationally expensive. Some studies (Chen et al., 2023f; Liu et al., 2023d) propose new optimizers to accelerate the pre-training of LLMs. Chen et al. (2023f) propose to leverage search techniques to traverse a large and sparse program space to discover optimizers for model training. The discovered optimizer, named Lion, is more memory-efficient than Adam as it only keeps track of the momentum. Liu et al. (2023d) propose Sophia as a lightweight second-order optimizer that outpaces Adam with doubling the pre-training speed. Sophia calculates the moving average of gradients and the estimated Hessian, dividing the former by the latter and applying element-wise clipping. It effectively moderates update sizes, addresses non-convexity and rapid hessian changes, enhancing both memory utilization and efficiency.
@@
System-Level Pre-Training Efficiency Optimization. Due to the high demand on memory and compute resources, LLMs are usually pre-trained across multiple compute nodes in a distributed manner. Therefore, most techniques for improving pre-training efficiency at the system level focus on distributed training. Existing efficient distributed training methods that are used for general AI model training can also be applied to LLM pre-training. For example, data parallelism (Li et al., 2020; Shallue et al., 2018) involves splitting the training dataset into multiple subsets on separate nodes. Each node computes gradients independently and then shares them with others to update the model parameters. Pipeline parallelism (Huang et al., 2019; Narayanan et al., 2019) divides the input minibatch into several smaller batches, and then distributes the execution of these microbatches across multiple GPUs. Tensor parallelism (Xu and You, 2023; Narayanan et al., 2021; Wang et al., 2022f; Bian et al., 2021) splits the model’s weight matrices across multiple nodes. Each node is responsible for executing the forward and backward passes with a segment of the model’s weights and their computed results are then aggregated. Although these parallelism techniques tackle the computing and memory constraints for training LLMs, they are still limited in maintaining computation, communication and development efficiency when fitting all of the runtime states, including gradients, optimizer states and activation states into limited memory. To bridge this gap, Zero Redundancy Data Parallelism (ZeRO) (Rajbhandari et al., 2020) provides three stages of optimization to partition the intermediate states during pre-training across different nodes. Specifically, ZeRO-1 only partitions the optimizer states, and ZeRO-2 partitions both the optimizer states and the gradients. Both ZeRO-1 and ZeRO-2 reduce runtime memory compared to data parallelism, while only consuming the same communication volume as data parallelism. ZeRO-3 provides a more aggressive partitioning that also splits the model parameter across the nodes compared with ZeRO-1 and ZeRO-2. Although runtime memory is further reduced through ZeRO-3, there is a modest 50% increase in communication overhead under this stage. Therefore, it is recommended to use ZeRO-3 within a node to minimize the communication time while using ZeRO-1 and ZeRO-2 across nodes. Fully Sharded Data Parallel (FSDP) (Zhao et al., 2023a) shares a similar idea for optimization, and designs a hybrid sharding strategy to allow users to define which nodes or processes to partition the gradients, parameter, and optimizer states across different nodes. In the case when the weight memory exceeds the aggregated memory that can be provided by all of the compute nodes, ZeRO-Offload (Ren et al., 2021) enables offloading to CPU for any stage of ZeRO, and ZeRO-Infinity (Rajbhandari et al., 2021a) provides a way to offload to NVMe drives in addition to CPU memory. However, it is quite difficult to maintain performance using these two alternatives, as the data movement between CPU and GPU is slow.
2.3. Efficient Fine-Tuning
@@
Efficient fine-tuning aims to enhance the efficiency of the fine-tuning process for LLMs. As shown in Figure 8, efficient fine-tuning methods can be grouped into parameter-efficient fine-tuning (PEFT), and memory-efficient fine-tuning (MEFT).
forked edges,
for tree=
grow=east,
reversed=true,
anchor=base west,
parent anchor=east,
child anchor=west,
base=center,
font=,
rectangle,
draw=hidden-draw,
rounded corners,
align=left,
text centered,
minimum width=4em,
edge+=darkgray, line width=1pt,
s sep=3pt,
inner xsep=2pt,
inner ysep=3pt,
line width=0.8pt,
where level=1text width=15em,font=,,
where level=2text width=12em,font=,,
Efficient Fine-Tuning, ver
Parameter-Efficient Fine-Tuning, fill=blue!10
Adapter-based Tuning, fill=blue!10
LLM-Adapters (Hu et al., 2023b), Compacter (Karimi Mahabadi et al., 2021), (IA) (Liu et al., 2022b),
Meta-Adapters (Bansal et al., 2022), AdaMix (Wang
et al., 2022d), OpenDelta (Hu
et al., 2023a), leaf, text width=23.5em
Low-Rank Adaptation, fill=blue!10
LoRA (Hu et al., 2022), LoRA-FA (Zhang
et al., 2023g), LoraHub (Huang
et al., 2023), LongLoRA (Chen
et al., 2023g),
Multi-Head Routing (Caccia et al., 2023), AdaLoRA (Zhang et al., 2023a),
Dylora (Valipour et al., 2022), CEPT (Zhao
et al., 2023b), Tied-LoRA (Renduchintala et al., 2023), leaf, text width=26em
Prefix Tuning, fill=blue!10
Prefix-Tuning (Li and Liang, 2021),
LLaMA-Adapter (Zhang et al., 2023b), leaf, text width=18em
Prompt Tuning, fill=blue!10
Soft Prompt (Lester
et al., 2021),
P-Tuning (Liu et al., 2021b),
P-Tuning v2 (Liu
et al., 2021a),
Tam et al. (2022),
MP (Sun
et al., 2022b),
PPT (Gu
et al., 2022b),
Multitask Prompt Tuning (Wang et al., 2023e), leaf, text width=23em
Memory-Efficient Fine-Tuning, fill=blue!10
QLoRA (Dettmers et al., 2023a),
QA-LoRA (Xu et al., 2023b),
LoftQ (Li et al., 2023f),
PEQA (Kim
et al., 2023b),
Selective Fine-Tuning (Simoulin
et al., 2023),
LOMO (Lv
et al., 2023),
MeZO (Malladi et al., 2023), leaf, text width=24em
2.3.1. Parameter-Efficient Fine-Tuning
Parameter-efficient fine-tuning (PEFT) aims to adapt an LLM to downstream tasks by freezing the whole LLM backbone and only updating a small set of extra parameters. In general, PEFT methods can be grouped into four categories: adapter-based tuning, low-rank adaptation, prefix tuning, and prompt tuning.
Adapter-based Tuning. Adapters are bottleneck-like trainable modules integrated into LLMs, which first down-project the input feature vector followed by a non-linear layer and then up-project back to the original size (Houlsby et al., 2019). Adapter-based tuning includes both series adapters and parallel adapters. In series adapters, each LLM layer has two adapter modules added after its attention and feed-forward modules; parallel adapters position two adapter modules alongside the attention and feed-forward modules within each layer of the LLM. In particular, Hu et al. (2023b) propose LLM-Adapters, which integrates series or parallel adapters into LLMs for fine-tuning on different tasks. Karimi Mahabadi et al. (2021) propose Compacter which unifies adapters, low-rank techniques, and the latest hyper-complex multiplication layers to achieve a balanced trade-off between the amount of trainable parameters and task performance. (IA) (Liu et al., 2022b) introduces a technique that scales activations using learned vectors, which outperforms few-shot in-context learning (ICL) in both accuracy and computational efficiency. Following meta-learning principles, Meta-Adapters (Bansal et al., 2022) designs a resource-efficient fine-tuning technique for the few-shot scenario where it incorporates adapter layers that have been meta-learned into a pre-trained model, transforming the fixed pre-trained model into an efficient few-shot learning framework. AdaMix (Wang et al., 2022d) takes inspiration from sparsely-activated mixture-of-experts (MoE) models (Zuo et al., 2021) and proposes a mixture of adaptation modules to learn multiple views of the given task. Lastly, OpenDelta (Hu et al., 2023a) is an open-source software library that offers a versatile and plug-and-play framework for implementing a range of adapter-based techniques, and is designed to be compatible with various LLMs architectures.
Low-Rank Adaptation. Low-Rank Adaptation (LoRA) (Hu et al., 2022) is a widely used PEFT approach for LLMs. Instead of directly adjusting the weight matrix as , LoRA introduces two trainable low-rank matrices and and expresses as . As such, only the small matrices and are updated during fine-tuning, while the original large weight matrix remains frozen, making the fine-tuning process more efficient. Though effective, LoRA still requires the update of all the parameters of the low-rank matrices for all the layers of the LLM at every single fine-tuning iteration. To enhance the efficiency of LoRA, LoRA-FA (Zhang et al., 2023g) keeps the projection-down weights of A fixed while updating the projection-up weights of B in each LoRA adapter so that the weight modifications during fine-tuning are confined to a low-rank space, thereby eliminating the need to store the full-rank input activations. LoraHub (Huang et al., 2023) explores the composability of LoRA for the purpose of generalizing across different tasks. It combines LoRA modules that have been trained on various tasks with the goal of attaining good performance on tasks that have not been seen before. LongLoRA (Chen et al., 2023g) extends LoRA to the long-context fine-tuning scenario. It introduces shift short attention (S-Attn), which effectively facilitates context expansion, showing that LoRA is effective for long context when utilizing trainable embedding and normalization. Multi-Head Routing (MHR) (Caccia et al., 2023) extends LoRA to Mixture-of-Experts (MoE) architectures. It outperforms Polytropon (Ponti et al., 2023) when operating with a similar parameter allocation. Notably, it achieves competitive performance while focusing on fine-tuning the routing function alone, without making adjustments to the adapters, demonstrating remarkable parameter efficiency. Zhang et al. (2023a) observe that many PEFT techniques neglect the differing significance of various weight parameters. To address this, they propose AdaLoRA which employs singular value decomposition to parameterize incremental updates and adaptively distributes the parameter budget based on the importance score of each weight matrix. Valipour et al. (2022) identify that the rank in LoRA is static and cannot be adaptively adjusted during fine-tuning. To address this issue, they propose Dylora, which introduces a dynamic low-rank adaptation method that trains LoRA blocks across multiple ranks rather than just one by organizing the representations learned by the adapter module based on their ranks. Different from above-mentioned methods that mainly apply PEFT to full-size LLMs, CEPT (Zhao et al., 2023b) introduces a new framework that utilizes compressed LLMs. Specifically, it assesses how prevalent LLM compression methods affect PEFT performance and subsequently implements strategies for knowledge retention and recovery to counteract the loss of knowledge induced by such compression techniques. Furthermore, Tied-LoRA (Renduchintala et al., 2023) uses weight tying and selective training to further increase parameter efficiency of LoRA.
Prefix Tuning. Prefix-Tuning (Li and Liang, 2021) adds a series of trainable vectors, known as prefix tokens, to each layer in an LLM. These prefix tokens are tailored to specific tasks and can be treated as virtual word embeddings. LLaMA-Adapter (Zhang et al., 2023b) incorporates a set of trainable adaptation embeddings and attaches them to the word embeddings in the upper layers of the LLMs. A zero-initialized attention scheme with zero gating is also introduced. It dynamically incorporates new guiding signals into LLaMA-1 while retaining its pre-trained knowledge.
Prompt Tuning. Different from prefix tuning, prompt tuning incorporates trainable prompt tokens at the input layer. These tokens can be inserted either as a prefix or anywhere within the input tokens. Soft Prompt (Lester et al., 2021) keeps the entire pre-trained model fixed while adding an extra trainable tokens at the beginning of the input text for each downstream task. It outperforms few-shot prompts and narrows the performance gap compared to full model fine-tuning. P-Tuning (Liu et al., 2021b) utilizes a small number of parameters as prompts, which are processed by a prompt encoder before being used as input for pre-trained LLMs. Instead of searching for discrete prompts, P-Tuning fine-tunes these prompts through gradient descent and improves performance on a wide range of NLU task. Liu et al. (2021a) observe that earlier versions of prefix tuning struggle with complex sequence labeling tasks. To address this, they propose P-Tuning v2 which enhances prefix tuning by introducing continuous prompts at each layer of the pre-trained model, rather than at the input layer only. This modification has proven effective in boosting performance across various parameter sizes for tasks related to natural language understanding. Tam et al. (2022) introduce efficient prompt tuning for text retrieval, updating just 0.1% of parameters and outperforming traditional full-parameter update methods in diverse domains. Sun et al. (2022b) claim that prompt tuning tends to struggle in few-shot learning scenarios, and thus propose MP that pre-trains a collection of modular prompts using multitask learning. These prompts are then selectively triggered and assembled by a trainable routing mechanism for specific tasks. As a result, MP can quickly adapt to downstream tasks by learning how to merge and reuse pretrained modular prompts. Different from MP, PPT (Gu et al., 2022b) attributes the performance degradation of prompt tuning in few-shot learning to the poor initialization of soft prompt, and thus proposes to add the soft prompt into the pre-training stage for a better initialization. Multitask Prompt Tuning (Wang et al., 2023e) harnesses the knowledge of the various tasks through the use of prompt vectors in a multitask learning setting. Specifically, it initially learns a single, transferable prompt by extracting knowledge from various task-specific source prompts, and then applies multiplicative low-rank updates to this prompt to effectively tailor it for each downstream task. By doing this, Multitask Prompt Tuning is able to attain performance levels that are competitive compared to full fine-tuning methods.
2.3.2. Memory-Efficient Fine-Tuning
As the parameters of LLMs expand, the sizes of memory needed for fine-tuning also increase, making memory a significant hurdle in fine-tuning. Consequently, minimizing memory usage in fine-tuning for improving efficiency has also emerged as a critical topic. Dettmers et al. (2023a) propose QLoRA which first quantizes the model into a 4-bit NormalFloat data type, and then fine-tunes this quantized model with added low-rank adapter (LoRA) weights (Hu et al., 2022). In doing so, QLoRA reduces memory usage during fine-tuning without performance degradation compared to standard full-model fine-tuning. QA-LoRA (Xu et al., 2023b) improves QLoRA by introducing group-wise operators that improve quantization flexibility (each group is quantized separately) while reducing adaptation parameters (each group utilizes shared adaptation parameters). Similarly, LoftQ (Li et al., 2023f) combines model quantization with singular value decomposition (SVD) to approximate the original high-precision pre-trained weights. As a result, it offers a favorable initialization point for subsequent LoRA fine-tuning, leading to enhancements over QLoRA. PEQA (Kim et al., 2023b) introduces a two-stage approach to quantization-aware fine-tuning. In the first stage, the parameter matrix for each fully connected layer is quantized into a matrix of low-bit integers along with a scalar vector. In the second stage, the low-bit matrix remains unchanged, while fine-tuning is focused solely on the scalar vector for each specific downstream task. Employing this two-stage approach, PEQA not only minimizes memory usage during fine-tuning but also speeds up inference time by maintaining weights in a low-bit quantized form. Simoulin et al. (2023) propose Selective Fine-Tuning which minimizes memory usage by specifically preserving a subset of intermediate activations from the forward pass for which the calculated gradients are nonzero. Notably, this approach delivers performance equivalent to full fine-tuning while using just up to one-third of the GPU memory required otherwise. Lv et al. (2023) introduce LOMO, which minimizes memory consumption during fine-tuning by combining gradient calculation and parameter updating into a single step. As such, LOMO eliminates all components of the optimizer state, lowering the memory requirements for gradient tensors to (1). MeZO (Malladi et al., 2023) improves the zeroth-order method (Spall, 1992) for gradient estimation using only two forward passes. This enables efficient fine-tuning of LLMs with memory requirements similar to inference and supports both full-parameter and PEFT methods like LoRA (Hu et al., 2022) and prefix tuning (Li and Liang, 2021), enabling MeZO to train a 30-billion parameter model on a single A100 80GB GPU.
@@ 2.3.2. Memory-Efficient Fine-Tuning
where level=2text width=18em,font=,,

### v4

@@ 2.2 Efficient Pre-Training
Training Optimizers. Popular LLMs such as GPT-3 (Brown et al., 2020), OPT (Zhang et al., 2022a), BLOOM (Scao et al., 2023), and Chinchilla (Hoffmann et al., 2022) are predominately pre-trained using Adam (Kingma & Ba, 2017) or AdamW (Loshchilov & Hutter, 2019) as optimizers. However, both Adam and AdamW are memory hungry and computationally expensive. Some studies (Chen et al., 2023g; Liu et al., 2024a) propose new optimizers to accelerate LLM pre-training. Specifically, Chen et al. (2023g) propose to leverage search techniques to traverse a large and sparse program space to discover optimizers for model training. The discovered optimizer, named Lion (EvoLved Sign Momentum), is more memory-efficient than Adam as it only keeps track of the momentum. Liu et al. (2024a), on the other hand, propose Sophia as a lightweight second-order optimizer that outpaces Adam with doubling the pre-training speed. Sophia calculates the moving average of gradients and the estimated Hessian, dividing the former by the latter and applying element-wise clipping. It effectively moderates update sizes, addresses non-convexity and rapid hessian changes, enhancing both memory utilization and efficiency.
@@
System-Level Pre-Training Efficiency Optimization. Due to high demand on memory and compute resources, LLMs are usually pre-trained across multiple compute nodes in a distributed manner. Therefore, most system-level optimization techniques are designed in the setting of large-scale distributed training. For instance, Zero Redundancy Data Parallelism (ZeRO) (Rajbhandari et al., 2020) provides three stages of optimization to partition various training states across different devices. Specifically, ZeRO-1 only partitions the optimizer states, whereas ZeRO-2 partitions both the optimizer states and the gradients. ZeRO-3 further partitions the model parameters across devices compared with ZeRO-1 and ZeRO-2. Although runtime memory is further reduced through ZeRO-3, there is about 50% increase in communication volume. Therefore, it is recommended to use ZeRO-3 within a node to minimize the communication time while using ZeRO-1 and ZeRO-2 across nodes. Fully Sharded Data Parallel (FSDP) (Zhao et al., 2023c) shares a similar idea for optimization, and designs a hybrid sharding strategy to allow users to define which nodes or processes to partition the gradients, model parameters, and optimizer states across different nodes. In the case when the weight memory exceeds the aggregated memory that can be provided by all of the compute nodes, ZeRO-Offload (Ren et al., 2021) enables offloading any stage of ZeRO to CPU memory, whereas ZeRO-Infinity (Rajbhandari et al., 2021) provides a mechanism to offload to NVMe drives in addition to CPU memory. However, it is quite difficult to maintain performance using these two alternatives, as the data movement between CPU and GPU is slow. Lastly, training LLMs on numerous GPUs consumes a massive amount of energy, Zeus (You et al., 2023) and Perseus (Chung et al., 2023) are proposed to optimize energy consumption by finding the best GPU-level configurations based on the unique LLM characteristics. The evaluation shows that Perseus reduces energy consumption of large model training by up to 30%.
2.3 Efficient Fine-Tuning
@@
Efficient fine-tuning techniques focus on reducing the costs of the LLM fine-tuning process. As summarized in Figure 8, efficient fine-tuning techniques can be grouped into parameter-efficient fine-tuning (PEFT) and memory-efficient fine-tuning (MEFT).
@@
where level=2text width=13em,font=,,
Efficient Fine-Tuning, ver
Parameter-Efficient Fine-Tuning, fill=blue!10
Low-Rank Adaptation, fill=blue!10
LoRA (Hu et al., 2022), LoRA-FA (Zhang et al., 2023b),
LoraHub (Huang et al., 2023), LongLoRA (Chen et al., 2023i),
Multi-Head Routing (Caccia et al., 2023), AdaLoRA (Zhang et al., 2023d),
DyLoRA (Valipour et al., 2023), CEPT (Zhao et al., 2023b),
Tied-LoRA (Renduchintala et al., 2023), leaf, text width=33em
Adapter-based Tuning, fill=blue!10
LLM-Adapters (Hu et al., 2023b),
Compacter (Karimi Mahabadi et al., 2021),
(IA)3 (Liu et al., 2022a),
Meta-Adapters (Bansal et al., 2022),
AdaMix (Wang et al., 2022c),
OpenDelta (Hu et al., 2023a),
SparseAdapter (He et al., 2022b), leaf, text width=34.5em
Prefix Tuning, fill=blue!10
Prefix-Tuning (Li & Liang, 2021),
LLaMA-Adapter (Zhang et al., 2024)
HyperTuning (Phang et al., 2023), leaf, text width=33em
Prompt Tuning, fill=blue!10
Prompt Tuning (Lester et al., 2021),
P-Tuning (Liu et al., 2023b),
P-Tuning v2 (Liu et al., 2022c),
Tam et al. (2023),
MP2 (Sun et al., 2023a),
PPT (Gu et al., 2022b),
Multitask Prompt Tuning (Wang et al., 2023j),
Xu et al. (2023b), leaf, text width=33.5em
Memory-Efficient Fine-Tuning, fill=blue!10
QLoRA (Dettmers et al., 2023),
QA-LoRA (Xu et al., 2024b),
LoftQ (Li et al., 2024d),
PEQA (Kim et al., 2023a),
Selective Fine-Tuning (Simoulin et al., 2023),
LOMO (Lv et al., 2023),
MeZO (Malladi et al., 2023),
Liu et al. (2023g), leaf, text width=51em
2.3.1 Parameter-Efficient Fine-Tuning
Parameter-efficient fine-tuning (PEFT) adapts an LLM to downstream tasks by freezing the whole LLM backbone and only updating a small set of newly added extra parameters. In general, PEFT methods can be grouped into four categories: low-rank adaptation, adapter-based tuning, prefix tuning, and prompt tuning.
Low-Rank Adaptation. Low-rank adaptation (LoRA) (Hu et al., 2022) is a widely used PEFT approach for LLMs. The hypothesis is that the change in weights during model adaptation has a low “intrinsic rank”. Hence, LoRA introduces two trainable low-rank matrices and and adjusts the weight matrix by . As such, only the small matrices and are updated during fine-tuning, while the original large weight matrix remains frozen, making the fine-tuning process more efficient. To enhance the efficiency of LoRA, LoRA-FA (Zhang et al., 2023b) keeps the projection-down weights of A fixed while only updating the projection-up weights of B in each LoRA adapter so that the weight modifications during fine-tuning are confined to a low-rank space, thereby eliminating the need to store the full-rank input activations. It achieves comparable accuracy related to full parameter fine-tuning and LoRA. Building on top of LoRA, LoraHub (Huang et al., 2023) explores the composability of LoRA for the purpose of generalizing across different tasks. It combines LoRA modules that have been trained on various tasks with the goal of attaining good performance on tasks that have not been seen before. LongLoRA (Chen et al., 2023i), on the other hand, extends LoRA to the long-context fine-tuning scenario. It introduces shift short attention (S2-Attn), which effectively facilitates context expansion, showing that LoRA is effective for long context when utilizing trainable embedding and normalization. Multi-Head Routing (MHR) (Caccia et al., 2023) extends LoRA to Mixture-of-Experts (MoE) architectures. It outperforms Polytropon (Ponti et al., 2023) when operating with a similar parameter allocation. Notably, it achieves competitive performance while focusing on fine-tuning the routing function alone, without making adjustments to the adapters, demonstrating remarkable parameter efficiency. Zhang et al. (2023d) observe that many PEFT techniques neglect the differing significance of various weight parameters. To address this, they propose AdaLoRA which employs singular value decomposition to parameterize incremental updates and adaptively distributes the parameter budget based on the importance score of each weight matrix. The rank in LoRA is static and cannot be adaptively adjusted during fine-tuning. Valipour et al. (2023) propose DyLoRA to introduce a dynamic low-rank adaptation method that trains LoRA blocks across multiple ranks rather than just one by organizing the representations learned by the adapter module based on their ranks. Different from the above-mentioned methods that apply LoRA-based methods to full-size LLMs, CEPT (Zhao et al., 2023b) introduces a framework that utilizes compressed LLMs. Specifically, it assesses how prevalent LLM compression methods affect PEFT performance and subsequently implements strategies for knowledge retention and recovery to counteract the loss of knowledge induced by compression. Lastly, Tied-LoRA (Renduchintala et al., 2023) uses weight tying and selective training to further increase parameter efficiency of LoRA.
Adapter-based Tuning. Adapters are bottleneck-like trainable modules integrated into LLMs, which first down-project the input feature vector followed by a non-linear layer and then up-project back to the original size (Houlsby et al., 2019). Adapter-based tuning includes both series adapters and parallel adapters. In series adapters, each LLM layer has two adapter modules added after its attention and feed-forward modules; whereas parallel adapters position two adapter modules alongside the attention and feed-forward modules within each layer of the LLM. In particular, Hu et al. (2023b) propose LLM-Adapters, which integrates series or parallel adapters into LLMs for fine-tuning on different tasks. Karimi Mahabadi et al. (2021) propose Compacter, which unifies adapters, low-rank techniques, and the latest hyper-complex multiplication layers to achieve a balanced trade-off between the amount of trainable parameters and task performance compared to original Adapter method (Houlsby et al., 2019). Furthermore, (IA)3 (Liu et al., 2022a) introduces a technique that scales activations using learned vectors. It outperforms Adapter (Houlsby et al., 2019) and Compacter on few-shot setting with better accuracy and computational efficiency. Following meta-learning principles, Meta-Adapters (Bansal et al., 2022) designs a resource-efficient fine-tuning technique for the few-shot scenario where it incorporates adapter layers that have been meta-learned into a pre-trained model, transforming the fixed pre-trained model into an efficient few-shot learning framework. Meta-Adapters outperforms Adapter (Houlsby et al., 2019) at few-shot fine-tuning with less parameters to fine-tune. AdaMix (Wang et al., 2022c) takes inspiration from sparsely-activated mixture-of-experts (MoE) models (Zuo et al., 2022) and proposes a mixture of adaptation modules to learn multiple views of the given task. Compared to Adapter, it demonstrates better results on both natural language understanding and generation tasks with less learnable parameters. Lastly, OpenDelta (Hu et al., 2023a) is an open-source software library that offers a versatile and plug-and-play framework for implementing a range of adapter-based techniques, and is designed to be compatible with various LLMs architectures.
Prefix Tuning. Prefix-Tuning (Li & Liang, 2021) adds a series of trainable vectors, known as prefix tokens, to each layer in an LLM. These prefix tokens are tailored to specific tasks and can be treated as virtual word embeddings. Building on top of Prefix-Tuning, LLaMA-Adapter (Zhang et al., 2024) incorporates a set of trainable adaptation embeddings and attaches them to the word embeddings in the upper layers of the LLMs. A zero-initialized attention scheme with zero gating is also introduced. It dynamically incorporates new guiding signals into LLaMA-1 while retaining its pre-trained knowledge. Different from conventional prefix tuning, HyperTuning (Phang et al., 2023) employs a hyper-model to produce task-specific parameters such as soft prefixes for a downstream model, showing improved performance through initialization from hypermodel-generated parameters for subsequent fine-tuning.
Prompt Tuning. Different from prefix tuning, prompt tuning incorporates trainable prompt tokens only at the input layer. These tokens can be inserted either as a prefix or anywhere within the input tokens. Prompt Tuning (Lester et al., 2021) keeps the entire pre-trained model fixed while adding an extra trainable tokens at the beginning of the input text for each downstream task. It outperforms few-shot prompts and narrows the performance gap compared to full-model fine-tuning. P-Tuning (Liu et al., 2023b) utilizes a small number of parameters as prompts, which are processed by a prompt encoder before being used as input for pre-trained LLMs. Instead of searching for discrete prompts, P-Tuning fine-tunes these prompts through gradient descent and improves performance on a wide range of natural language understanding tasks compared to Prompt Tuning. Liu et al. (2022c) observe that earlier versions of prefix tuning struggle with complex sequence labeling tasks. To address this, they propose P-Tuning v2, which borrows the ideas from prefix tuning by introducing continuous prompts at each layer of the pre-trained model. This modification has proven effective in boosting performance across various parameter sizes for tasks related to natural language understanding. Tam et al. (2023) introduce efficient prompt tuning for text retrieval, updating just 0.1% of parameters and outperforming traditional full-parameter update methods in diverse domains. Sun et al. (2023a) claim that prompt tuning tends to struggle in few-shot learning scenarios, and thus propose MP2 that pre-trains a collection of modular prompts using multitask learning. These prompts are then selectively triggered and assembled by a trainable routing mechanism for specific tasks. As a result, MP2 can quickly adapt to downstream tasks by learning how to merge and reuse pretrained modular prompts. Different from MP2, PPT (Gu et al., 2022b) attributes the performance degradation of prompt tuning in few-shot learning to the poor initialization of soft prompt, and thus proposes to add the soft prompt into the pre-training stage for a better initialization. Lastly, Multitask Prompt Tuning (Wang et al., 2023j) extends Prompt Tuning and harnesses the knowledge of the various tasks through the use of prompt vectors in a multitask learning settings. Specifically, it initially learns a single, transferable prompt by extracting knowledge from various task-specific source prompts, and then applies multiplicative low-rank updates to this prompt to effectively tailor it for each downstream task. By doing this, Multitask Prompt Tuning is able to attain performance levels that are competitive compared to full-model fine-tuning methods.
2.3.2 Memory-Efficient Fine-Tuning
Different from PEFT methods which focus on parameter efficiency, MEFT methods focus on memory savings during the LLMs fine-tuning process. For instance, Dettmers et al. (2023) propose QLoRA which first quantizes the model into a 4-bit NormalFloat data type, and then fine-tunes this quantized model with added low-rank adapter (LoRA) weights (Hu et al., 2022). In doing so, QLoRA reduces memory usage during fine-tuning without performance degradation compared to standard full-model fine-tuning. QA-LoRA (Xu et al., 2024b) improves QLoRA by introducing group-wise operators that improve quantization flexibility (each group is quantized separately) while reducing adaptation parameters (each group utilizes shared adaptation parameters). Similarly, LoftQ (Li et al., 2024d) combines model quantization with singular value decomposition (SVD) to approximate the original high-precision pre-trained weights. As a result, it offers a favorable initialization point for subsequent LoRA fine-tuning, leading to enhancements over QLoRA on both natural language understanding and generation tasks. PEQA (Kim et al., 2023a) introduces a two-stage approach to quantization-aware fine-tuning. In the first stage, the parameter matrix for each fully connected layer is quantized into a matrix of low-bit integers along with a scalar vector. In the second stage, the low-bit matrix remains unchanged, while fine-tuning is focused solely on the scalar vector for each specific downstream task. Employing this two-stage approach, PEQA not only minimizes memory usage during fine-tuning but also speeds up inference time by maintaining weights in a low-bit quantized form, showing better perplexity than GPTQ (Frantar et al., 2023) with LoRA. Different from above-mentioned MEFT methods that combine LoRA with quantization to reduce fine-tuning memory footprints, as another line of research, some studies propose MEFT methods based on gradient optimization. Specifically, Simoulin et al. (2023) propose Selective Fine-Tuning which minimizes memory usage by specifically preserving a subset of intermediate activations from the forward pass for which the calculated gradients are nonzero. Notably, this approach delivers performance equivalent to full-model fine-tuning while using just up to one-third of the GPU memory required otherwise. Lv et al. (2023) introduce LOMO, which minimizes memory consumption during fine-tuning by combining gradient calculation and parameter updating into a single step. As such, LOMO eliminates all components of the optimizer state, lowering the memory requirements for gradient tensors to (1). Lastly, MeZO (Malladi et al., 2023) improves the zeroth-order method (Spall, 1992) for gradient estimation using only two forward passes. This enables efficient fine-tuning of LLMs with memory requirements similar to inference and supports both full-parameter and PEFT methods like LoRA (Hu et al., 2022) and prefix tuning (Li & Liang, 2021), enabling MeZO to train a 30-billion parameter model on a single A100 80GB GPU.
2.4 Efficient Inference
Efficient inference techniques focus on reducing the costs of the LLMs inference process. As summarized in Figure 10, efficient inference techniques can be grouped into techniques at algorithm level and system level.
forked edges,
for tree=
grow=east,
reversed=true,
anchor=base west,
parent anchor=east,
child anchor=west,
base=center,
font=,
rectangle,
draw=hidden-draw,
rounded corners,
align=left,
text centered,
minimum width=4em,
edge+=darkgray, line width=1pt,
s sep=3pt,
inner xsep=2pt,
inner ysep=3pt,
line width=0.8pt,
where level=1text width=16em,font=,,
where level=2text width=17em,font=,,